A violin string vibrates with a frequency of 440 Hz in the fundamental mode. With what frequency will the string vibrate if 5 loops are observed?
____________Hz
Answer #1

As we nkow that

f = nf0

Here

f0 = Fundamental Frequency = 440 Hz

Therefore

f' = 5*440

= 2200 Hz
Answer #2

A violin is a both ends fixed, and so successive harmonics are simply multiples of the fundamental:
Therefore, f1 = 440Hz (fundamental frequency)
f2 = 2f1 = 2 x 440 = 880 Hz
f3 = 3f1 = 3 x 440 = 1320 Hz
f4 = 4f1 = 4 x 440 = 1760 Hz
f5 = 5f1 = 5 x 440 = 2200 Hz
Answer #3

Fundamental frequency = V/2L =440 Hz

For 5 loops

Frequency = 5V/2L =2200 Hz (ANS)
Answer #4

Frequency=n*Fundamental frequency

=5*440

=1760 Hz ans